OpenStudy (anonymous):

Solve the equation. (x + 1)2 = 9

OpenStudy (acxbox22):

(x+1)2=9 x+1=4.5 -1 -1 x=3.5

OpenStudy (anonymous):

how did you get the 4.5

OpenStudy (anonymous):

2 (x+1) = 9